# GimbalDiffusion: Gravity-Aware Camera Control for Video Generation

### v1

@@ 5 Discussion
While our method introduces a novel capability for text-to-video models, it has a few limitations. First, our current on-the-fly camera sampling supports only rotational motion. Extending to translations—potentially by leveraging recent successes in real-time novel view synthesis methods such as Gaussian Splats—remains an interesting direction for future work. Second, the generated videos can exhibit visual artifacts, as visible in the generated crowd in fig. 6. We expect that continued progress in video generation models will mitigate these issues while remaining compatible with our framework.
6 Conclusion
We presented GimbalDiffusion, the first method to enable absolute camera control in text-to-video generation by using gravity as global reference. Our data pipeline combines 360° panoramic videos with camera pose estimation to produce gravity-aligned camera pose annotations. Leveraging panoramic videos allows us to sample the full sphere of possible viewpoints, reducing the domain gap for rare but expressive camera angles. In addition, we proposed null-pitch conditioning, a data generation step that enhances camera conditioning fidelity when the model is given prompts with conflicting cues.
We believe our work lays the foundation for physically grounded camera control in video generation, empowering creators to produce dynamic storyboards and convey their visual ideas with greater precision and speed.

### v2

@@ 5 Discussion
We present GimbalDiffusion, a method that enables extreme and arbitrary camera control in text-to-video generation by grounding camera poses in a gravity-aligned absolute coordinate system. Our pipeline combines 360° panoramic videos with camera pose estimation to produce training data covering the full sphere of possible viewpoints, including extreme pitch, large roll, and full 360° rotation trajectories that were previously beyond the reach of generative video models. We also proposed null-pitch conditioning, a data generation step that enhances camera conditioning fidelity when the text prompts contain conflicting cues.
While it introduces novel capabilities for text-to-video models, our method still has a few limitations. First, our current on-the-fly camera sampling is restricted to rotations and must use the translations from the original source videos. A promising area for future research is to extend this data augmentation to include translations, potentially by leveraging recent successes in real-time novel view synthesis methods such as 3DGS [21]. Second, the generated videos can exhibit visual artifacts (for example, the faceless man in fig. 4). We expect that continued progress in video generation models will mitigate these issues while remaining compatible with our framework. We hope our work lays the foundation for robust camera control under extreme rotations in video generation, empowering creators to produce dynamic storyboards with arbitrary camera orientations and convey their visual ideas with greater precision and speed.